# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v1

@@ 5.2 Improved Optimistic Value Function Construction
In Section 5.1, we improve the dependence on $\kappa$ by constructing a tighter $\kappa$-independent confidence set. However, the regret of UCRL-MNL-LL in Theorem 2 still have a dependence on $\kappa^{-1/2}$. In this part, we design an improved algorithm that removes the dependence on $\kappa$ in the main term.
By a similar analysis as (7), the regret of UCRL-MNL-LL can be bounded as:
$\operatorname*{Reg}(K)\leq 2H\widetilde{\beta}_{K}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+\sum_{k=1}^{K}\sum_{h=1}^{H}\mathcal{M}_{k,h}.$ (11)
Though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (11), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence. To address this issue, we propose a new construction of the optimistic value function through a more refined analysis. Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from the estimation error in the following lemma.
@@
Suppose Lemma 3 holds. For any function $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, with probability at least $1-\delta$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}},$
@@
$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\quad\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
@@
$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (12)
The detailed algorithm is presented in UCRL-MNL-LL+ in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL+ in the following theorem.
@@
Suppose Assumption 1 holds. For any $\delta\in(0,1)$, set $\eta=\frac{1}{2}\log(1+U)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$ and $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, with probability at least $1-\delta$, UCRL-MNL-LL+ algorithm ensures the following regret guarantee
@@
Theorem 3 moves the exponentially small $\kappa$-dependence to the lower-order term, which is independent of $K$. Compared with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023), the higher-order term only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.
@@ 6 Lower Bound
In this section, we establish the first lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and logistic bandits.
Consider the following logistic bandit problem (Faury et al., 2020): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d}\mid\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter. The learner aims to minimize the regret: ${\operatorname*{Reg}}^{\mathtt{LB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*})$.
Theorem 4.
For any logistic bandit $\mathcal{B}$, there exists an MNL mixture MDP $\mathcal{M}$ such that learning $\mathcal{M}$ is as hard as learning $H/2$ independent instances of $\mathcal{B}$ simultaneously.
@@
Corollary 1 can be proved by combining Theorem 4 and the $\Omega(d\sqrt{K\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al. (2021). To our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work. Consequently, it remains open for the lower bound of MNL mixture MDPs with finite action space.
Remark 5.
While Corollary 1 is not a strict lower bound for MNL mixture MDPs with finite action space, it gives a basis to conjecture our results may be optimal in $d$ and $K$.
Dependence on $H$.
MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021). The pioneering work by Ayoub et al. (2020) achieved an $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ regret for linear mixture MDPs, which matches our results in Theorem 3, differing only on the lower-order term. Zhou et al. (2021) enhanced the dependence on $H$ and attained an optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$. This was made possible by recognizing that the value function in linear mixture MDPs is also linear, allowing for direct learning of value functions while incorporating their variance information. In contrast, the value function does not conform to a logistic model for MNL mixture MDPs, posing a significant challenge in using variance information of value functions. It remains open whether similar improvements in $H$ are attainable for MNL mixture MDPs.
Dependence on $\kappa$.
Though we fully remove the $\kappa$ dependence in Theorem 3 in the higher-order term and move it to the lower order term, Corollary 1 suggests that there may still be some room for improvement in the dependency on $\kappa$. The $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}})$ dependence has been established in the binary logistic bandit problem by Abeille et al. (2021). However, their techniques are limited to the binary case and are not directly applicable to MNL mixture MDPs which have multinomial rewards. For the multinomial feedback settings, Périvier and Goyal (2022) studied the MNL contextual bandit problem and achieved a regret bound of $\widetilde{\mathcal{O}}(d\sqrt{K\kappa^{*}}+\kappa^{-1})$. However, their results only hold for the scenarios with uniform revenue, where each choice yields identical rewards. In contrast, MNL mixture MDPs are more challenging as they involve diverse revenue values, represented by the value functions of different states. Thus, their techniques also can not be applied to MNL mixture MDPs. It remains open whether the $\widetilde{\mathcal{O}}(dH\sqrt{K\kappa^{*}})$ result can be achieved for MNL mixture MDPs.
@@ 7 Conclusion and Future Work
In this work, we address both the computational and statistical challenges for MDPs with multinomial logit function approximation. We first propose an algorithm that attains the previously best-known $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. Next, we identify that the statistical efficiency can be further improved by exploiting the local information. We propose two algorithms that not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve improved regrets of $\widetilde{\mathcal{O}}(\kappa^{-1/2}dH^{2}\sqrt{K})$ and $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, respectively. We also establish the first lower bound, justifying the optimality of our results in $d$ and $K$. To the best of our knowledge, this is the first work that achieves almost the same computational and statistical efficiency as linear function approximation while employing non-linear function approximation.
There are several interesting directions for future work. First, there still exists a gap between the upper and lower bounds and how to close this gap is an open problem. Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Fei et al., 2020; Li et al., 2023) is another interesting direction.
@@ Appendix A Properties of the Multinomial Logit Function
This section collects several key properties of the multinomial logit function.
@@
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),\quad g_{k,h}(\theta)=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
@@
$(1)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
$(2)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(3)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
$(4)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(5)\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
@@
We prove the three statements individually.
Proof of statement (1).
By the definition that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$, we have
@@
$\leq\frac{4}{\kappa}\sum_{i=1}^{k}\log\Bigg{(}1+\min\bigg{\{}1,\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}\bigg{\}}\Bigg{)}$
@@
where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by Equation (13).
Proof of statement (2).
By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Thus, we have
@@
Proof of statement (3).
By Lemma 5, we have $\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\leq\kappa$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$. By a similar analysis as the statement (2), we have
@@
Proof of statement (4).
By the definition of $H_{k,h}(\theta)$, we have
@@
Proof of statement (5).
By (14), we have

### v2

@@ 5.2 Improved Optimistic Value Function Construction
Based on the confidence set in Lemma 3, a natural choice for the optimistic value function construction is analogous to (7) and can be expressed as:
$\bar{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\bar{V}_{k,h+1}(s^{\prime})+H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\|\phi_{s,a}^{s^{\prime}}\|_{\mathcal{H}_{k,h}^{-1}}\bigg{]}_{[0,H]}.$ (10)
However, though $\widetilde{\beta}_{K}$ now is independent of $\kappa$ and $\mathcal{H}_{k,h}^{-1}$ preserves the local information in (10), the norm $\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is still in a global manner due to the maximum operation, leading to a $\kappa^{-1/2}$ dependence. To address this issue, we propose a new construction of the optimistic value function. Specifically, we employ a second-order Taylor expansion to more accurately bound the value difference arising from transition estimation errors.
Although the idea of using second-order Taylor expansion has been explored in bandits (Périvier and Goyal, 2022; Zhang and Sugiyama, 2023; Lee and Oh, 2024), fundamental differences arise in the MDP setting. Specifically, Périvier and Goyal (2022) studied the uniform revenue setting, where the reward is identical for all actions. Zhang and Sugiyama (2023) and Lee and Oh (2024) focused on the non-uniform setting, but the rewards for different actions are known a priori. However, in the MDP setting, the value function is state-dependent and unknown to the learner, leading to a more challenging problem. Moreover, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward. Due to these differences, the analyses used in the bandit setting cannot be directly applied to the MDP setting.
For MDPs, we show the value difference arising from the transition estimation error as follows.
@@
Suppose Lemma 3 holds. For any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}.$
@@
$\epsilon_{s,a}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{s,a}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
@@
$\widetilde{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widetilde{V}_{k,h+1}(s^{\prime})+\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}\bigg{]}_{[0,H]}.$ (11)
In contrast to the value function specified in (10), where the term $\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ is utilized, the refined value function introduced in (11) substitutes this term with $\epsilon_{s,a}^{\mathtt{fst}}+\epsilon_{s,a}^{\mathtt{snd}}$. This adjustment better preserves local information, offering a more precise and $\kappa$-independent estimation error bound.
At state $s_{k,h}$, our algorithm chooses action $a_{k,h}=\max\widetilde{Q}_{k,h}(s_{k,h},\cdot)$. The detailed algorithm is presented in Algorithm 2. We show the regret guarantee of UCRL-MNL-LL in the following theorem.
@@
For any $\delta\in(0,1)$, set $\widetilde{\beta}_{k}=\mathcal{O}(\sqrt{d}\log U\log(KH/\delta))$, $\eta=\frac{1}{2}\log(1+U)+(B+1)$ and $\lambda=84\sqrt{2}\eta(B+d)$, with probability at least $1-\delta$, UCRL-MNL-LL algorithm ensures the following regret guarantee
@@
The high-order term in Theorem 2 is now independent $\kappa$, significantly improving the statistical efficiency compared with Theorem 1. In comparison with the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear cases (Zhou et al., 2021; He et al., 2023), the higher-order term in Theorem 2 only differs by a factor of $H^{1/2}$, almost matching the same computational and statistical efficiency simultaneously.
@@ 6 Lower Bound
In this section, we establish a lower bound for MNL mixture MDPs by presenting a novel reduction, which connects MNL mixture MDPs and the logistic bandit problem.
Consider the following logistic bandit problem (Faury et al., 2020; Abeille et al., 2021): at each round $k\in[K]$, the learner selects an action $x_{k}\in\mathcal{X}$ and receives a reward $r_{k}$ sampled from Bernoulli distribution with mean $\mu(x^{\top}\theta^{*})=(1+\exp(-x^{\top}\theta^{*}))^{-1}$, where $\theta^{*}\in\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$ is the unknown parameter. The learner aims to to minimize the regret:
${\operatorname*{Reg}}^{\mathtt{LogB}}(K)=\max_{x\in\mathcal{X}}\sum_{k=1}^{K}\mu(x^{\top}\theta^{*})-\sum_{k=1}^{K}\mu(x_{k}^{\top}\theta^{*}).$
Theorem 3 (Lower Bound).
For any logistic bandit problem $\mathcal{B}$, there exists an MNL mixture MDP $\mathcal{M}$ such that learning $\mathcal{M}$ is as hard as learning $H/2$ independent instances of $\mathcal{B}$ simultaneously.
@@
Corollary 1 can be proved by combining Theorem 3 and the $\Omega(d\sqrt{K\kappa^{*}})$ lower bound for logistic bandits with infinite arms by Abeille et al. (2021). To the best of our knowledge, a lower bound for logistic bandits with finite arms has not been established, which is beyond the scope of this work. This absence leaves the lower bound for MNL mixture MDPs with a finite action space open through this reduction. However, after the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al. (2024) proposed a new reduction that bridges MNL mixture MDPs with linear mixture MDPs by approximating multinomial logit functions to linear functions, employing the mean value theorem. Leveraging this new reduction, they established an $\Omega(dH^{3/2}\sqrt{K})$ lower bound for MNL mixture MDPs with the finite action space. This achievement confirms that our result is indeed optimal with respect to the dependence on $d$ and $K$, only differing by a $\mathcal{O}(H^{1/2})$ factor.
Dependence on $H$.  By the discussion in Remark 7, we note that our result is optimal with respect to $d$ and $K$, but loosing by a $\mathcal{O}(H^{1/2})$ factor. We discuss the challenges in improving the dependence on $H$. Notably, MNL mixture MDPs can be viewed as a generalization of linear mixture MDPs (Ayoub et al., 2020; Zhou et al., 2021). The pioneering work by Ayoub et al. (2020) achieved a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K})$ for linear mixture MDPs, which matches our results in Theorem 2, differing only on the lower order term. Later, Zhou et al. (2021) enhanced the dependence on $H$ and attained an optimal regret bound of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$. This was made possible by recognizing that the value function in linear mixture MDPs is linear, allowing for direct learning of the value function while incorporating variance information. In contrast, the value function for MNL mixture MDPs does not conform to a specific structure, posing a significant challenge in using the variance information of value functions. Thus, it remains open whether similar improvements on $H$ are attainable for MNL mixture MDPs.
@@ 7 Conclusion and Future Work
In this work, we study MNL mixture MDPs that employ multinomial logit function approximation to ensure valid probability distributions over the state space. We address both the computational and statistical challenges for this problem. Specifically, we first propose an algorithm based on the online Newton step that attains the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. Next, we propose an enhanced algorithm that leverages local information to improve the statistical efficiency. It not only maintains $\mathcal{O}(1)$ computational and storage costs but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly matching the result of linear function approximation from both computational and statistical perspectives. Finally, we establish the first lower bound for MNL mixture MDPs, justifying the optimality of our results in $d$ and $K$.
There are several interesting directions for future work. First, there still exists a gap between the upper and lower bounds and how to close this gap is an open problem. Besides, while this work focuses on stochastic rewards, extending this model to the non-stationary settings and studying the dynamic regret (Wei and Luo, 2021; Zhao et al., 2022; Li et al., 2023) is an another important direction.
Acknowledgments
This research was supported by National Science and Technology Major Project (2022ZD0114800) and NSFC (U23A20382, 62206125). Peng Zhao was supported in part by the Xiaomi Foundation.
@@ Appendix A Properties of the Multinomial Logit Function
This section collects several key properties of the multinomial logit function used in the paper.
@@
$f_{k,h}(\theta)$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}y_{k,h}^{s^{\prime}}\log p_{k,h}^{s^{\prime}}(\theta),$
$g_{k,h}(\theta)$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$
@@
$\widehat{\Sigma}_{k,h}$ $=\frac{\kappa}{2}\sum_{i=1}^{k-1}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}+\lambda I_{d},$
$\mathcal{H}_{k,h}$ $=\sum_{i=1}^{k-1}H_{i,h}(\widetilde{\theta}_{i+1,h})+\lambda I_{d},$
@@
$(\textnormal{I})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i,h}^{-1}}^{2}\leq\frac{4}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)$
(II$)\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{k,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)( I$
$(\textnormal{III})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
$(\textnormal{IV})\quad\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq 2d\log\left(1+\frac{k}{\lambda d}\right)$
$(\textnormal{V})\quad\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\widetilde{\phi}_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right)$
@@
We prove the five statements individually.
Proof of statement (I).  By the definition that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$, we have
@@
where the first inequality holds by the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$ and the last inequality holds by (12).
Proof of statement (II).  By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Thus, we have
@@
Proof of statement (III).  By Lemma 5, we have
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\leq\kappa$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$. By a similar analysis as the statement (2), we have
@@
This finishes the proof.
Proof of statement (IV).  By the definition of $H_{k,h}(\theta)$, we have
@@
Proof of statement (V).  By (13), we have